OpenStudy (anonymous):

how can I find the derivative of the following function. S(t)=t+(9/(t+1))+1

OpenStudy (anonymous):

ds/dt as dy/dx

OpenStudy (anonymous):

the only rule I really understand for derivatives is the power rule and I don't know how to apply it to fractions

OpenStudy (anonymous):

\[ds/dt=t+\frac{ 9 }{ t+1}+1\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

here you have 3 parts

OpenStudy (anonymous):

ds/dt=t ds/dt=(9/(t+1)) ds/dt=1

OpenStudy (anonymous):

i think you know the dreiv. of the 1st one and the last one

OpenStudy (anonymous):

the derivative of the first is 1 and there isn't one for the last one right?

OpenStudy (anonymous):

the second one it has his one rule

OpenStudy (anonymous):

yes about the first one. would you mean by there isn't one for the last one? do mean the answer is (0) or there isn't one in the equation

OpenStudy (anonymous):

the answer is 0

OpenStudy (anonymous):

I honestly thought it meant that there wasn't one present, but I guess you are implying that if there is not a variable then the derivative=0

OpenStudy (anonymous):

df/dx= for a constant =0

OpenStudy (anonymous):

2nd (f'*g-fg')/g^2

OpenStudy (anonymous):

this the rule f=9 g=t+1

OpenStudy (anonymous):

so my answer would be 1-(9/(t+1)^2)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

awesome! thanks. Also, how would I solve that equation for t?

OpenStudy (anonymous):

never mind I see it. thank you